# Nonlinearity-induced symmetry breaking in a system of two parametrically driven Kerr-Duffing oscillators

### v2

@@ Appendix A The model Hamiltonian
with $q_{n}$ and $p_{n}$ the conjugated variables (i.e. position and momentum) of the two modes. The dynamics equations are given by
@@
$\tilde{H}_{n}$ $=\frac{1}{4}{\left(Q_{n}^{2}+P_{n}^{2}\right)}^{2}-\frac{\delta\Omega_{n}}{2}{\left(Q_{n}^{2}+P_{n}^{2}\right)}$
$+\frac{\delta\Omega_{T,n}}{2}{\left(Q_{n}^{2}-P_{n}^{2}\right)}$ (38)
$\tilde{H}_{int}$ $={\left(Q_{1}^{2}+P_{1}^{2}\right)}{\left(Q_{2}^{2}+P_{2}^{2}\right)}+\frac{1}{2}{\left(Q_{1}^{2}-P_{1}^{2}\right)}{\left(Q_{2}^{2}-P_{2}^{2}\right)}$
$+2Q_{1}Q_{2}P_{1}P_{2}$ (39)
@@
$\frac{1}{\Gamma_{1}}\frac{\partial z_{1}}{\partial t}$ $=\left(-i\delta\Omega_{1}+i{|z_{1}|}^{2}-1\right)z_{1}+i\delta\Omega_{T,1}z_{1}^{*}$
$+i\frac{g}{\alpha}\left(2{|z_{2}|}^{2}z_{1}+z_{2}^{2}z_{1}^{*}\right)$ (41)
$\frac{1}{\Gamma_{2}}\frac{\partial z_{2}}{\partial t}$ $=\left(-i\delta\Omega_{2}+i{|z_{2}|}^{2}-1\right)z_{2}+i\delta\Omega_{T,2}z_{2}^{*}$
$+i\alpha g\left(2{|z_{1}|}^{2}z_{2}+z_{2}^{1}z_{2}^{*}\right)\,.$ (42)
@@ Appendix B The noninteracting case $g=0$ and the steady-state trivial solution $\bar{z}_{1}=\bar{z}_{2}=0$
$\bar{z}^{(0)}_{1,+}=\pm i\sqrt{\delta\Omega_{1}+\delta\Omega_{T,1}}\,,\,\,\,\,\bar{z}^{(0)}_{1,-}=\pm\sqrt{\delta\Omega_{1}-\delta\Omega_{T,1}}\,,$ (44)
The solution $\bar{z}^{(0)}_{1,+}$ corresponds to the the resonator state when it starts to oscillate at frequency $\omega_{d}$ above the detuning threshold $\delta\Omega>-\delta\Omega_{T,1}$ and it is completely out of phase with respect to the drive. By analyzing the dynamics of the harmonic fluctuations around the stationary point, this solution is always dynamically stable above the threshold. The solution $\bar{z}^{(0)}_{1,-}$ appears above the detuning threshold $\delta\Omega>\delta\Omega_{T,1}$ and it is in phase respect to the drive: This solution is always dynamically unstable above the corresponding threshold. The trivial solution $\bar{z}^{(0)}_{1}=0$ is unstable in the detuning range $|\delta\Omega|<\delta\Omega_{T,1}$.
The effect of finite damping $\Gamma\neq 0$ is a shift of the detuning threshold $\delta\Omega_{T,1}\rightarrow{(\delta\Omega_{T,1}^{2}-1)}^{1/2}$, which reads, using the bare parameters, as the change from $\mu_{1}/(2\omega_{d})\rightarrow{[{(\mu_{1}/(2\omega_{d}))}^{2}-\Gamma^{2}]}^{1/2}$.
@@ Appendix C Solutions with $\bar{z}_{2}=0$
However, the two resonators are correlated as the finite value of the amplitude of the first resonator $\bar{z}_{1}$ affects the fluctuations $\delta z_{2}$ of the second one with $\bar{z}_{2}=0$. Therefore, the stability range can differ from that of the noninteracting case. As a consequence, the stationary pair solution, which is stable for the case $g=0$ $\left(\bar{z}_{1,+}=\pm i\sqrt{\delta\Omega+\delta\Omega_{T,1}},\bar{z}_{2}=0\right)$ can become unstable. For example, in the regime $\omega_{1}=\omega_{2}$, $\delta\Omega_{T,2}<g\delta\Omega_{T,1}$ and $g<1/2$, the zero amplitude solution for the second resonator $\bar{z}_{2}=0$ becomes unstable in the detuning range $\delta\Omega_{T,2}<\delta\Omega<\left(2g\delta\Omega_{T,1}-\delta\Omega_{T,2}\right)/(1-2g)$, well beyond the range of the noninteracting case $|\delta\Omega|<\delta\Omega_{T,2}$.
@@ Appendix D Analytic solutions for $\delta\Omega_{T,1}>0$ and $\delta\Omega_{T,2}=0$
$0=\left({|z_{1}|}^{2}-\delta\Omega+2g{|z_{1}|}^{2}\right)z_{2}+\left(z_{1}^{2}\right)z_{2}^{*}$ (45)
@@
The solutions in table 1 are possible solutions. However, as second step, we analyze the dynamical stability of these stationary solutions. In general, the region of stability does not coincide with the detuning threshold at which the stationary solution appears, and we must hence compute the stability ranges numerically.
@@
Condition Range Discontinuity $\bar{z}_{1}$ $\bar{z}_{2}$
$g<1$ $\delta\Omega>-\delta\Omega_{T,1}$ $\delta\Omega_{c,1}=\frac{g}{1-g}\delta\Omega_{T,1}$$\pm i\sqrt{\delta\Omega+\delta\Omega_{T,1}-\theta\left(\delta\Omega-\delta\Omega_{c,1}\right)\frac{g}{1+g}\left(\delta\Omega-\delta\Omega_{c,1}\right)}$ $\mp\sqrt{\delta\Omega-g{|\bar{z}_{1}|}^{2}}$
$g<1$ $\delta\Omega>\delta\Omega_{c,2}$ $\delta\Omega_{c,2}=\frac{1}{1-g}\delta\Omega_{T,1}$$\pm\sqrt{\frac{1}{1+g}\left(\delta\Omega-\delta\Omega_{c,2}\right)}$$\pm i\sqrt{\delta\Omega-g{|\bar{z}_{1}|}^{2}}$
$1<g$ $\delta\Omega>\delta\Omega_{c,1}^{*}$ $\delta\Omega_{c,1}^{*}=\frac{1}{g-1}\delta\Omega_{T,1}$$\pm i\sqrt{\frac{1}{1+g}\left(\delta\Omega-\delta\Omega_{c,1}^{*}\right)}$ $\mp\sqrt{\delta\Omega-g{|\bar{z}_{1}|}^{2}}$
$1<g$ $\delta\Omega>\delta\Omega_{T,1}$ $\delta\Omega_{c,2}^{*}=\frac{g}{g-1}\delta\Omega_{T,1}$$\pm\sqrt{\delta\Omega-\delta\Omega_{T,1}-\theta\left(\delta\Omega-\delta\Omega_{c,2}^{*}\right)\frac{g}{1+g}\left(\delta\Omega-\delta\Omega_{c,2}^{*}\right)}$$\pm i\sqrt{\delta\Omega-g{|\bar{z}_{1}|}^{2}}$
$g<\frac{1}{3}$ $\delta\Omega>-\delta\Omega_{T,1}$ $\delta\Omega_{c,3}=\frac{3g}{1-3g}\delta\Omega_{T,1}$$\pm i\sqrt{\delta\Omega+\delta\Omega_{T,1}-\theta\left(\delta\Omega-\delta\Omega_{c,3}\right)\frac{3g}{1+3g}\left(\delta\Omega-\delta\Omega_{c,3}\right)}$$\pm i\sqrt{\delta\Omega-3g{|\bar{z}_{1}|}^{2}}$
$g<\frac{1}{3}$ $\delta\Omega>\delta\Omega_{c,4}$ $\delta\Omega_{c,4}=\frac{1}{1-3g}\delta\Omega_{T,1}$$\pm\sqrt{\frac{1}{1+3g}\left(\delta\Omega-\delta\Omega_{c,4}\right)}$$\pm\sqrt{\delta\Omega-3g{|\bar{z}_{1}|}^{2}}$
$\frac{1}{3}<g$ $\delta\Omega>\delta\Omega_{c,3}^{*}$ $\delta\Omega_{c,3}^{*}=\frac{1}{3g-1}\delta\Omega_{T,1}$$\pm i\sqrt{\frac{1}{1+3g}\left(\delta\Omega-\delta\Omega_{c,3}^{*}\right)}$$\pm i\sqrt{\delta\Omega-3g{|\bar{z}_{1}|}^{2}}$
$\frac{1}{3}<g$ $\delta\Omega>\delta\Omega_{T,1}$ $\delta\Omega_{c,4}^{*}=\frac{3g}{3g-1}\delta\Omega_{T,1}$$\pm\sqrt{\delta\Omega-\delta\Omega_{T,1}-\theta\left(\delta\Omega-\delta\Omega_{c,4}^{*}\right)\frac{3g}{1+3g}\left(\delta\Omega-\delta\Omega_{c,4}^{*}\right)}$$\pm\sqrt{\delta\Omega-3g{|\bar{z}_{1}|}^{2}}$
Table 1: Non-trivial analytic solutions of the stationary state for a single driven mode, for the symmetric case, and in the limit
small damping.

### v3

@@ Appendix A The model Hamiltonian
with $q_{n}$ and $p_{n}$ the conjugated variables (i.e. position and momentum) of the two modes. The dynamical equations are given by
@@
$\tilde{H}_{n}$ $=\frac{1}{4}{\left(Q_{n}^{2}+P_{n}^{2}\right)}^{2}-\frac{\delta\Omega_{n}}{2}{\left(Q_{n}^{2}+P_{n}^{2}\right)}+\frac{\delta\Omega_{T,n}}{2}{\left(Q_{n}^{2}-P_{n}^{2}\right)}\,,$ (38)
$\tilde{H}_{int}$ $={\left(Q_{1}^{2}+P_{1}^{2}\right)}{\left(Q_{2}^{2}+P_{2}^{2}\right)}+\frac{1}{2}{\left(Q_{1}^{2}-P_{1}^{2}\right)}{\left(Q_{2}^{2}-P_{2}^{2}\right)}+2Q_{1}Q_{2}P_{1}P_{2}\,,\phantom{AAAA}$ (39)
@@
$\frac{1}{\Gamma_{1}}\frac{\partial z_{1}}{\partial t}$ $=\left(-i\delta\Omega_{1}+i{|z_{1}|}^{2}-1\right)z_{1}+i\delta\Omega_{T,1}z_{1}^{*}+i\frac{g}{\alpha}\left(2{|z_{2}|}^{2}z_{1}+z_{2}^{2}z_{1}^{*}\right)$ (41)
$\frac{1}{\Gamma_{2}}\frac{\partial z_{2}}{\partial t}$ $=\left(-i\delta\Omega_{2}+i{|z_{2}|}^{2}-1\right)z_{2}+i\delta\Omega_{T,2}z_{2}^{*}+i\alpha g\left(2{|z_{1}|}^{2}z_{2}+z_{2}^{1}z_{2}^{*}\right)\,.$ (42)
@@ Appendix B The noninteracting case $g=0$ and the steady-state trivial solution $\bar{z}_{1}=\bar{z}_{2}=0$
$\bar{z}^{(0)}_{1,+}=\pm i\sqrt{\delta\Omega_{1}+\delta\Omega_{T,1}}\,,\,\,\,\,\bar{z}^{(0)}_{1,-}=\pm\sqrt{\delta\Omega_{1}-\delta\Omega_{T,1}}\,.$ (44)
The solution $\bar{z}^{(0)}_{1,+}$ corresponds to the resonator state when it starts to oscillate at frequency $\omega_{d}$ above the detuning threshold $\delta\Omega>-\delta\Omega_{T,1}$ and it is completely out of phase with respect to the drive. By analyzing the dynamics of the harmonic fluctuations around the stationary point, this solution is always dynamically stable above the threshold. The solution $\bar{z}^{(0)}_{1,-}$ appears above the detuning threshold $\delta\Omega>\delta\Omega_{T,1}$ and it is in phase respect to the drive: This solution is always dynamically unstable above the corresponding threshold. The trivial solution $\bar{z}^{(0)}_{1}=0$ is unstable in the detuning range $|\delta\Omega|<\delta\Omega_{T,1}$.
For finite damping the scaled equation reads
$\left({|z_{1}|}^{2}-\delta\Omega_{1}+i\right)z_{1}=-\delta\Omega_{T,1}z_{1}^{*}\,,$ (45)
and taking the square modulus of the left and right side of Eq. (45), assuming a nontrivial solution $|z_{1}|\neq 0$, we have ${\left({|z_{1}|}^{2}-\delta\Omega_{1}\right)}^{2}=\delta\Omega_{T,1}^{2}-1$. This implies that the effect of finite damping $\Gamma\neq 0$ is a shift of the detuning threshold $\delta\Omega_{T,1}\rightarrow{(\delta\Omega_{T,1}^{2}-1)}^{1/2}$, which reads, using the bare parameters, as the change from $\mu_{1}/(2\omega_{d})\rightarrow{[{(\mu_{1}/(2\omega_{d}))}^{2}-\Gamma^{2}]}^{1/2}$.
@@ Appendix C Solutions with $\bar{z}_{2}=0$
However, the two resonators are correlated as the finite value of the amplitude of the first resonator $\bar{z}_{1}$ affects the fluctuations $\delta z_{2}$ of the second one with $\bar{z}_{2}=0$. Therefore, the stability range can differ from that of the noninteracting case. As a consequence, the stationary pair solution, which is stable for the case $g=0$ $\left(\bar{z}_{1,+}=\pm i\sqrt{\delta\Omega+\delta\Omega_{T,1}},\bar{z}_{2}=0\right)$ can become unstable. For example, in the regime $\omega_{1}=\omega_{2}$, $\delta\Omega_{T,2}<g\delta\Omega_{T,1}$ and $g<1/2$, the zero amplitude solution with the second resonator $\bar{z}_{2}=0$ becomes unstable in the detuning range $\delta\Omega_{T,2}<\delta\Omega<\left(2g\delta\Omega_{T,1}-\delta\Omega_{T,2}\right)/(1-2g)$, well beyond the range of the noninteracting case $|\delta\Omega|<\delta\Omega_{T,2}$.
@@ Appendix D Analytic solutions for $\delta\Omega_{T,1}>0$ and $\delta\Omega_{T,2}=0$
$0=\left({|z_{1}|}^{2}-\delta\Omega+2g{|z_{1}|}^{2}\right)z_{2}+\left(z_{1}^{2}\right)z_{2}^{*}\,.$ (46)
@@
Table 1 only reports stationary, possible solutions. However, as second step, we must analyze the dynamical stability of these stationary solutions. In general, the region of stability does not coincide with the detuning threshold at which the stationary solution appears, and we must hence compute the stability ranges numerically.
@@
Cond. Range Discontinuity $\bar{z}_{1}$ $\bar{z}_{2}$
$g<1$ $\delta\Omega>-\delta\Omega_{T,1}$ $\delta\Omega_{c,1}=\frac{g}{1-g}\delta\Omega_{T,1}$$\pm i\sqrt{\delta\Omega+\delta\Omega_{T,1}-\theta\left(\delta\Omega-\delta\Omega_{c,1}\right)\frac{g}{1+g}\left(\delta\Omega-\delta\Omega_{c,1}\right)}$ $\mp|\bar{z}_{2,+}|$
$g<1$ $\delta\Omega>\delta\Omega_{c,2}$ $\delta\Omega_{c,2}=\frac{1}{1-g}\delta\Omega_{T,1}$$\pm\sqrt{\frac{1}{1+g}\left(\delta\Omega-\delta\Omega_{c,2}\right)}$$\pm i|\bar{z}_{2,+}|$
$1<g$ $\delta\Omega>\delta\Omega_{c,1}^{*}$ $\delta\Omega_{c,1}^{*}=\frac{1}{g-1}\delta\Omega_{T,1}$$\pm i\sqrt{\frac{1}{1+g}\left(\delta\Omega-\delta\Omega_{c,1}^{*}\right)}$ $\mp|\bar{z}_{2,+}|$
$1<g$ $\delta\Omega>\delta\Omega_{T,1}$ $\delta\Omega_{c,2}^{*}=\frac{g}{g-1}\delta\Omega_{T,1}$$\pm\sqrt{\delta\Omega-\delta\Omega_{T,1}-\theta\left(\delta\Omega-\delta\Omega_{c,2}^{*}\right)\frac{g}{1+g}\left(\delta\Omega-\delta\Omega_{c,2}^{*}\right)}$$\pm i|\bar{z}_{2,+}|$
$g<\frac{1}{3}$ $\delta\Omega>-\delta\Omega_{T,1}$ $\delta\Omega_{c,3}=\frac{3g}{1-3g}\delta\Omega_{T,1}$$\pm i\sqrt{\delta\Omega+\delta\Omega_{T,1}-\theta\left(\delta\Omega-\delta\Omega_{c,3}\right)\frac{3g}{1+3g}\left(\delta\Omega-\delta\Omega_{c,3}\right)}$$\pm i|\bar{z}_{2,-}|$
$g<\frac{1}{3}$ $\delta\Omega>\delta\Omega_{c,4}$ $\delta\Omega_{c,4}=\frac{1}{1-3g}\delta\Omega_{T,1}$$\pm\sqrt{\frac{1}{1+3g}\left(\delta\Omega-\delta\Omega_{c,4}\right)}$$\pm|\bar{z}_{2,-}|$
$\frac{1}{3}<g$ $\delta\Omega>\delta\Omega_{c,3}^{*}$ $\delta\Omega_{c,3}^{*}=\frac{1}{3g-1}\delta\Omega_{T,1}$$\pm i\sqrt{\frac{1}{1+3g}\left(\delta\Omega-\delta\Omega_{c,3}^{*}\right)}$$\pm i|\bar{z}_{2,-}|$
$\frac{1}{3}<g$ $\delta\Omega>\delta\Omega_{T,1}$ $\delta\Omega_{c,4}^{*}=\frac{3g}{3g-1}\delta\Omega_{T,1}$$\pm\sqrt{\delta\Omega-\delta\Omega_{T,1}-\theta\left(\delta\Omega-\delta\Omega_{c,4}^{*}\right)\frac{3g}{1+3g}\left(\delta\Omega-\delta\Omega_{c,4}^{*}\right)}$$\pm|\bar{z}_{2,-}|$
Table 1: Non-trivial analytic solutions of the stationary state for a single driven mode, for the symmetric case, and in the limit of vanishing damping. The solution for $\bar{z}_{2}$ are $|\bar{z}_{2,+}|=\sqrt{\delta\Omega-g{|\bar{z}_{1}|}^{2}}$ and $|\bar{z}_{2,-}|=\sqrt{\delta\Omega-3g{|\bar{z}_{1}|}^{2}}$ (see text).